OpenStudy (anonymous):

9 times a number has 1 subtracted from it. the answer is trebled. the ans is the same if u add 7 to twice the num and multiply the result by 7. what is the number

OpenStudy (anonymous):

9x-1=y (7+2x)7=y 49+14x=9x-1 50=-5x x=-10

OpenStudy (anonymous):

thanks